OpenStudy (unofficialllyy):

Calculate the average speed (in km/h) of Larry, who runs to the store 4.6 kilometers away in 24 minutes.

OpenStudy (unofficialllyy):

I know that Average Speed= total distance covered/time interval So I do 4.6/24 and get 0.1916666667

OpenStudy (phi):

yes, but that is km/min

OpenStudy (unofficialllyy):

So how do I do km/h?

OpenStudy (phi):

60 min / 1 hour do this: \[0.1916666667 \frac{km}{\cancel{min}} \cdot \frac{60 \ \cancel{min}}{hour} = ? \]

OpenStudy (unofficialllyy):

Okay I cross multiplied and got 11.5

OpenStudy (phi):

or you could first change 24 minutes to hours doing \[ 24 \ \cancel{min} \cdot \frac{1 \ hour}{60\ \cancel{min}} = 0.4\ hours\] then do \[ \frac{4.6 \ km}{0.4 \ hrs} \]

OpenStudy (unofficialllyy):

Right, but when I put in 11.5 km/h it says its wrong. it says "Not quite. Check through your calculations; you may have made a rounding error or used the wrong number of significant figures."

OpenStudy (phi):

maybe it just wants 11.50 ?

OpenStudy (unofficialllyy):

Nope. :/

OpenStudy (phi):

the only other thing I can think of, is it want only 2 significant digits, which would mean 12

OpenStudy (unofficialllyy):

There's a space for value which is where I put 11.5 and then there's a space for unit which is where I put km/h

OpenStudy (phi):

can you post a screen shot ?

OpenStudy (unofficialllyy):

Yes one second

OpenStudy (phi):

btw, did you try 12 km/h ?

OpenStudy (unofficialllyy):

I didn't even think to round up, yes it worked lol. Thanks!